OpenStudy (destinyyyy):

Perform the addition (-9/2 +1/2i) + (3/2 - 7/2i)

OpenStudy (destinyyyy):

My example shows to put the i on the outside of the equation like this- (-9/2 +1/2) + (3/2 - 7/2)i Im not sure what to do next

OpenStudy (usukidoll):

I wouldn't do it like that. Our answer has to be in the form of ai+b or a+bi... the same rule applies for subtraction ai-b or a-bi so what I would've done is to combine the like terms as in put all the terms without the i's to the left and all terms with the i's to the right

OpenStudy (usukidoll):

\[\frac{-9}{2}+\frac{3}{2}+\frac{1}{2}i-\frac{7}{2}i\]

OpenStudy (usukidoll):

now the denominators are the same which is good because we don't have to use GCF or LCD ... so can you tell me what -9+3 is? and 1-7 ?

OpenStudy (destinyyyy):

-6

OpenStudy (usukidoll):

alright so now we have \[\frac{-6}{2}+\frac{-6}{2}i\] so what's -6 divided by 2

OpenStudy (destinyyyy):

Im trying to follow but Id like to follow the examples with the i on the outside

OpenStudy (usukidoll):

you need to rearrange the terms a bit before you can factor the i outside.

OpenStudy (destinyyyy):

I did another problem like this-- (-5+5i) + (9+8i) = (-5+5) + (9+8)i -5+9 + 5+8 4 + 13i

OpenStudy (usukidoll):

there's a problem with that

OpenStudy (destinyyyy):

What would that be?? The answer is correct

OpenStudy (usukidoll):

(-5+5i)+(9+8i) you have to rearrange the terms first it should look something like this -5+9+5i+8i and then it would be 4+13i which is right, but you need to do some rearranging first

OpenStudy (usukidoll):

either combine like terms or take the i outside -5+9+(5+8)i 4+13i

OpenStudy (usukidoll):

the same goes with this \[\frac{-9}{2}+\frac{3}{2}+\frac{1}{2}i-\frac{7}{2}i \] \[\frac{-9}{2}+\frac{3}{2}+(\frac{1}{2}-\frac{7}{2})i\]

OpenStudy (usukidoll):

O_O! What happened to my latex? -____-! OS is acting up.

OpenStudy (destinyyyy):

All of my examples do it with the i on the outside so its easier for me to understand

OpenStudy (destinyyyy):

latex?

OpenStudy (usukidoll):

there's more than one method to doing this problem and it all gives the same result as long as you do the correct arithmetic.

OpenStudy (destinyyyy):

Okay

OpenStudy (destinyyyy):

Im following you now

OpenStudy (usukidoll):

let me rewrite the top portion... it's not displaying right

OpenStudy (destinyyyy):

Okay

OpenStudy (usukidoll):

\[\frac{-9}{2}+\frac{1}{2}i+\frac{3}{2}-\frac{7}{2}i\] \[\frac{-9}{2}+\frac{3}{2}+\frac{1}{2}i-\frac{7}{2}i \] [rearrange the terms] \[\frac{-9}{2}+\frac{3}{2}+(\frac{1}{2}-\frac{7}{2})i \] [factor the i so it appears on the outside] \[\frac{-6}{2}+(\frac{-6}{2})i \] [solve for the numerators] last step is division

OpenStudy (usukidoll):

-6/2 = ?

OpenStudy (destinyyyy):

-3

OpenStudy (usukidoll):

\[-3-3i \] yup and we're done :)

OpenStudy (destinyyyy):

Okay thank you!